Suppose that in Germany computers cost $1,500 and automobiles cost $20,000, and in the U.S. computers cost $1,000 and automobiles cost $15,000. Which of the following is not correct? 1. the U.S. has an absolute advantage in both computers and automobiles 2. all of the other answers are correct 3. Germany has a comparative advantage in automobiles 4. the U.S. will not gain if it trades with Germany 5. the U.S. has a comparative advantage in computers
Answer #1

Ans is 4.

both nations benefits after trading

Opportunity cost of computer in Germany=1500/2000=0.75automobiles

opportunity cost of computer in U.S.=1000/1500=0.667automobiles

thus US has comparative advantage in Computers and Germany in automobiles